OpenStudy (anonymous):

A carpenter makes legs for 3-legged stools and 4-legged tables. He makes a total of 46 legs. He could have made legs for how many 3-legged stools? A. 8 B. 10 C. 12 D. 16 help!!!

OpenStudy (anonymous):

If he is only making stools, he would only be able to make 15 stools, because 3 legs for each stool, 46 total legs, 46/3, correct?

OpenStudy (anonymous):

3s + 4t = 46 Sub in your answer choices in for s (A) 3(8) + 4t = 46 24 + 4t = 46 4t = 46 - 24 4t = 22 t = 22/4 = 5 1/2 This one will not work because your answer has to be a whole number with no fractions. Now try the 2nd one... (B) 3(10) + 4t = 46 30 + 4t = 46 4t = 46 - 30 4t = 16 t = 4 BINGO...this one works. 10 three legged stools and 4 four legged tables can be made. (C) 3(12) + 4t = 46 36 + 4t = 46 4t = 46 - 36 4t = 10 t = 10/4 = 2 1/2 (will not work) (D) 3(16) + 4t = 46 48 + 4t = 46 4t = 46 - 48 t = - 2 (will not work) any questions ?

OpenStudy (anonymous):

I don't understand how you arrived at using the equation you did (3s + 4t = 46) given the information they are asking for. They aren't asking to factor in tables, only stools.

OpenStudy (anonymous):

It would seem then that the answers are incorrect, no?

OpenStudy (anonymous):

@timo86m ...can you take a look at this. I thought I did the problem correctly, however, joshua has a point. They are only asking for stool legs. I am having doubts. What do you think ?

OpenStudy (anonymous):

@thomaster ...please take a look at this